# Optimization and Generalization Guarantees for Weight Normalization

### v1

@@ 1 Introduction
Weight normalization (WeightNorm) was first introduced by Salimans and Kingma [2016]. The idea is to normalize each weight vector–by dividing it by its Euclidean norm–in order to decouple its length from its direction. Salimans and Kingma [2016] reported that WeightNorm is able to speed-up the training of neural networks, as an alternative to the then recently introduced and still widely-used batch normalization (BatchNorm) [Ioffe and Szegedy, 2015]. WeightNorm is different from BatchNorm because it is independent from the statistics of the batch sample being used at the gradient step. It is also different from another class of normalization, called layer normalization (LayerNorm) [Ba et al., 2016] in that it does not entail the normalization of any activation function of the neural network. Thus WeightNorm also has less computational overhead because it does not require the computation and storage of additional mean and standard deviation statistics as the other methods do. However, its empirical testing performance (or accuracy) has often been found to not be as good as other normalization methods by itself, and thus various versions have been formulated to improve its performance [Salimans and Kingma, 2016, Huang et al., 2017, Qiao et al., 2020]. Nevertheless, WeightNorm has been one of the most popular normalization methods since its introduction, and as a result current machine learning libraries include built-in implementations of it, e.g., PyTorch 2. .
@@
Parallel to the empirical development of normalization methods, recent years have seen advances in theoretically understanding convergence of gradient descent (GD) and variants for deep learning models [Du et al., 2019, Allen-Zhu et al., 2019, Zou et al., 2020, Nguyen, 2021, Liu et al., 2022, Banerjee et al., 2023]. However, to the best of our knowledge, no prior work has focused on providing formal optimization and generalization guarantees for deep models with WeightNorm. In this paper, our contribution is to study the effect of WeightNorm on both training and generalization from a theoretical perspective. In particular, we focus on smooth networks–deep models with smooth activations–a setting which has become increasingly popular in recent years [Du et al., 2019, Huang and Yau, 2020, Liu et al., 2022, Banerjee et al., 2023].
[…]
@@
Our third technical contribution is to provide generalization guarantees to smooth networks with WeightNorm. We iteratively use WeightNorm properties and our network scaling to obtain a generalization gap which is independent from the width–our bound is ${\cal O}(\sqrt{L}/\sqrt{n})$, where $L$ is the depth and $n$ the sample size. Thus, we avoid an exponential dependence on the depth found in the literature of networks without WeightNorm [Bartlett et al., 2017, Neyshabur et al., 2018, Golowich et al., 2018]. Our generalization guarantee uses a uniform convergence approach based on Rademacher complexity.
[…]
@@ 2 Related work
WeightNorm was first proposed by Salimans and Kingma [2016], after a year from the introduction of BatchNorm by Ioffe and Szegedy [2015]. LayerNorm was proposed in the same year by Ba et al. [2016], and has also become very popular [Xu et al., 2019]. Extensions and variations of WeightNorm have been introduced since then, for example, centered weight normalization [Huang et al., 2017], orthogonal weight normalization [Huang et al., 2018], and weight standardization [Qiao et al., 2020]. Theoretical analysis of normalization techniques in general has caught an increasing recent interest. We focus mostly on WeightNorm, but mention that BatchNorm has some theoretical work, e.g. [Santurkar et al., 2018, Lian and Liu, 2019]. The work [Arpit et al., 2019] uses a theoretical approach to better come up with initialization schemes tailored for WeightNorm networks with ReLU activations. In contrast, our work focuses in understanding the optimization (training) and generalization of WeightNorm networks with smooth activations. Wu et al. [2020] study the effect of WeightNorm on the convergence of overparameterized least squares problem (which translates to a normalization of the parameters to estimate). We, in contrast, study the non-convex problem of optimizing WeighNorm networks. Dukler et al. [2020] present exponential convergence guarantees on shallow ReLU networks with WeightNorm, whereas we study smooth deep networks and present generalization too. We conclude by remarking that the field of normalization is growing in proposing new normalization schemes and improving their understanding and application–we refer the reader to the recent survey [Huang et al., 2023].
@@
Given the increasingly vast literature, we refer the readers to the surveys [Fan et al., 2021, Bartlett et al., 2021] for an overview of the field of gradient descent training of neural networks. Deep ReLU networks were analyzed by Zou and Gu [2019], Zou et al. [2020], Allen-Zhu et al. [2019], Nguyen and Mondelli [2020], Nguyen [2021], Nguyen et al. [2021], whereas deep networks with smooth activations–our case–were analyzed by Du et al. [2019], Liu et al. [2022]. Regarding the convergence analysis of gradient descent, Du et al. [2019], Allen-Zhu et al. [2019], Zou and Gu [2019], Zou et al. [2020], Liu et al. [2022] used the Neural Tangent Kernel (NTK) approach [Jacot et al., 2018]; whereas Banerjee et al. [2023] used the RSC approach. All training guarantees hold over a neighborhood around the initialization point of the network’s weights.
@@
To the best of our knowledge, the first work in studying the Rademacher complexity for neural networks was [Neyshabur et al., 2015]. Then, the seminal work [Bartlett et al., 2017] provided generalization bounds for neural networks considering Lipschitz activation functions. The bound, obtained using a covering approach, depends exponentially on the depth of the network. The generalization bound was tightened for ReLU activations in [Neyshabur et al., 2018] using a PAC-Bayes approach, but the same exponential dependency persisted. Then, the work [Golowich et al., 2018] avoided some exponential dependence on the depth and established conditions on the norm of the network’s weight matrices in order to obtain generalization bounds independent of depth. Unlike our work, all the cited works do not study WeightNorm networks–indeed, we show that WeightNorm allows the generalization bound to depend on $\sqrt{L}$, where $L$ is the depth. Finally, we remark that our generalization bound estimates how well finding a solution to the empirical loss approximates finding a solution to the loss under the real data distribution. A different problem in the literature is inductive bias, i.e., how a specific optimization method, while training the network, may converge to a solution that generalizes well. This was studied for WeightNorm under GD in [Morwani and Ramaswamy, 2022] and under stochastic GD in [Li et al., 2022].
@@ 3 Problem setup: deep learning with WeightNorm
${\left\lVert\bar{W}\right\rVert_{2}}={\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i}\right\rVert_{2}^{2}}$.
@@
Finally, we point out that the original WeightNorm paper by Salimans and Kingma [2016] also has a scalar scaling multiplying the normalized weight vectors $W_{i}^{(l)}/\left\lVert W_{i}^{(l)}\right\rVert_{2}$, $l\in[L]$. However, their work considers optimizing over such scaling, whereas we consider a fixed scaling for our analysis.
@@ 4 Characterizing WeightNorm bounds on the loss and predictor
[…]
@@
[…]
@@
Surprisingly, the Lipschitz constant of the network (7)–unlike networks without WeightNorm–does not explicitly depend on the depth $L$ of the network nor its width $m$. It depends on the radius $\rho_{1}$ which can be set to be a constant ${\cal O}(1)$. This is a direct effect of WeightNorm: when computing $\nabla_{\mathbf{x}}f(\theta;\mathbf{x}_{i})$, we use a chain-rule argument that depends on the product of Jacobians of the output $\alpha^{(l)}$ at layer $l$ with respect to the previous output $\alpha^{(l-1)}$, and we show each Jacobian has at most unit norm due to WeightNorm. Thus, no matter how deep the network is, the final effect this product of Jacobians has on the Lipschitz constant of the network is just a constant one. ∎
@@ 5 Optimization guarantees for WeightNorm
In this section we prove our training guarantees for WeightNorm networks under the square loss. We first introduce two technical lemmas that will be used for our convergence analysis, defining the restricted strong convexity (RSC) property and a smoothness-like property respectively. These two properties use the bounds derived in Section 4. All proofs are found in Section C of the appendix.
@@
$\begin{split}{\cal L}(\theta^{\prime})&\geq{\cal L}(\theta)+\langle\theta^{\prime}-\theta,\nabla_{\theta}{\cal L}(\theta)\rangle+\frac{\alpha_{\theta}}{2}\|\theta^{\prime}-\theta\|_{2}^{2}~{},\end{split}$ (12)
@@
$\alpha_{\theta}=\frac{\kappa^{2}}{2}\frac{\left\lVert\nabla_{\theta}{\cal L}(\theta)\right\rVert_{2}^{2}}{{\cal L}(\theta)}$ (13)
$-{\cal O}\left(\left(1+\frac{\sqrt{L}}{{\left\lVert\bar{W}\right\rVert_{2}}}\right)\frac{(1+\rho_{2})(1+\rho_{1})^{3}L^{3}A(\frac{1}{\sqrt{m}},L^{2},|\phi(0)|)B(1,L,|\phi(0)|\sqrt{m})}{\min\left\{{\left\lVert\bar{W}\right\rVert_{2}},{\left\lVert\bar{W}\right\rVert_{2}^{2}}\right\}}\right)$
where $A(\frac{1}{\sqrt{m}},L^{2},|\phi(0)|)=\frac{1}{\sqrt{m}}+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\}$ and $B(1,L,|\phi(0)|\sqrt{m})=1+L|\phi(0)|\sqrt{m}$. We say that the empirical loss ${\cal L}$ satisfies the RSC property w.r.t. $(Q^{\theta}_{\kappa}\cap B_{\rho_{2}}^{\operatorname{Euc}}(\theta),\theta)$ whenever $\alpha_{\theta}>0$.
[…]
@@
$\alpha_{\theta}=\frac{\kappa^{2}}{2}\frac{\left\lVert\nabla_{\theta}{\cal L}(\theta)\right\rVert_{2}^{2}}{{\cal L}(\theta)}-{\cal O}\left(\left(1+\frac{\sqrt{L}}{\sqrt{m}{\left\lVert\bar{W}\right\rVert_{2}}}\right)\frac{(1+\rho_{2})(1+\rho_{1})^{3}L^{3}}{\sqrt{m}\min\left\{{\left\lVert\bar{W}\right\rVert_{2}},{\left\lVert\bar{W}\right\rVert_{2}^{2}}\right\}}\right).$ (14)
[…]
@@
Our RSC characterization in the left-hand side of (13) depends on $\frac{\|\nabla_{\theta}\mathcal{L}(\theta)\|_{2}^{2}}{\mathcal{L}(\theta)}$, whereas [Banerjee et al., 2023]–the work that introduced the RSC-based optimization analysis–depends on $\|\frac{1}{n}\sum_{i=1}^{n}\nabla_{\theta}f(\theta,x_{i})\|_{2}^{2}$. Moreover, the minimum weight vector norm appears in the right-hand side since we used the bounds on Section 4. ∎
@@
[…]
@@
[…]
@@
${\cal L}(\theta^{\prime})\leq{\cal L}(\theta)+\langle\theta^{\prime}-\theta,\nabla_{\theta}{\cal L}(\theta)\rangle+\frac{\beta_{\theta}}{2}\|\theta^{\prime}-\theta\|_{2}^{2}~{},$ (15)
@@
$\beta_{\theta}\leq{\cal O}\left(L^{4}A(1,L^{2},|\phi(0)|)(1+\rho_{1})^{2}\left(1+\frac{1}{\min\{{\left\lVert\bar{W}\right\rVert_{2}},{\left\lVert\bar{W}\right\rVert_{2}^{2}}\}}\right)\right)~{},$ (16)
where $A(1,L^{2},|\phi(0)|)=(1+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\}m)^{2}$.
[…]
@@
$\beta_{\theta}\leq{\cal O}\left(1+\frac{L^{3}(1+\rho_{1})^{3}}{\sqrt{m}\min\{{\left\lVert\bar{W}\right\rVert_{2}},{\left\lVert\bar{W}\right\rVert_{2}^{2}}\}})\right).$ (17)
[…]
Remark 5.4 (The $\alpha_{\theta}/\beta_{\theta}$ ratio).
If the expressions in equations (12) and (15) hold for any $\theta,\theta^{\prime}\in\mathbb{R}^{p}$ so that $\alpha_{\theta}\equiv\alpha>0$ and $\beta_{\theta}\equiv\beta>0$, then we have the definitions of strong convexity and smoothness, respectively–moreover, it immediately follows that $\frac{\alpha}{\beta}<1$. Having such ratio between the parameters being less than the one is, in general, not guaranteed for the RSC and the smoothness-like properties of any function; however, it is in our case. From the proof of Lemma 5.2 it holds that $2\varrho^{2}_{\theta}<\beta_{\theta}$. Thus, using Proposition 4.1, we have that $\alpha_{\theta}<\frac{\kappa^{2}}{2}\frac{\left\lVert\nabla_{\theta}{\cal L}(\theta)\right\rVert_{2}^{2}}{{\cal L}(\theta)}\leq 2\varrho^{2}_{\theta}<\beta_{\theta}$, since $\kappa^{2}<1$, which implies $\frac{\alpha_{\theta}}{\beta_{\theta}}<1$.∎
The following assumptions are useful for establishing our convergence analysis–indeed, as we shortly show, a decrease on the loss value is guaranteed as long as these assumptions are satisfied.
@@
Consider iterates $\{\theta_{t}\}_{t=0,1,\dots,T}$ and that: (A3.1) $\alpha_{\theta_{t}}>0$; (A3.2) gradient descent (GD) update $\theta_{t+1}=\theta_{t}-\eta_{t}\nabla{\cal L}(\theta_{t})$ with learning rate $\eta_{t}$ chosen appropriately so that $\mathbf{v}_{t}\in B_{\rho_{1}}^{\operatorname{Euc}}(\mathbf{v}_{0})$; (A3.3) $\rho_{2}$ is chosen so that $\theta_{t+1}\in B_{\rho_{2}}^{\operatorname{Euc}}(\theta_{t})$.
The first statement in Assumption 3 ensures that the empirical loss satisfies the RSC property. The second statement ensures that gradient descent has an appropriate learning rate. Finally, the third statement defines a ball around the current iterate that should cover the next iterate–in principle, its radius could be defined after defining the learning rate of GD of the second assumption.
Finally, we present our main optimization result: a reduction on the empirical loss towards its minimum value (with the last layer in the Euclidean set of radius $\rho_{1}$) using gradient descent. This is proved by using Lemmas 5.1 and 5.2 and an adaptation of [Banerjee et al., 2023, Theorem 5.3].

### v2

@@ 1 Introduction
Weight normalization (WeightNorm) was first introduced by Salimans and Kingma [2016]. The idea is to normalize each weight vector—by dividing it by its Euclidean norm—in order to decouple its length from its direction. Salimans and Kingma [2016] reported that WeightNorm is able to speed-up the training of neural networks, as an alternative to the then recently introduced and still widely-used batch normalization (BatchNorm) [Ioffe and Szegedy, 2015]. WeightNorm is different from BatchNorm because it is independent from the statistics of the batch sample being used at the gradient step. It is also different from another class of normalization, called layer normalization (LayerNorm) [Ba et al., 2016] in that it does not entail the normalization of any activation function of the neural network. Thus WeightNorm also has less computational overhead because it does not require the computation and storage of additional mean and standard deviation statistics as the other methods do. However, its empirical testing performance (or accuracy) has often been found to not be as good as other normalization methods by itself, and thus various versions have been formulated to improve its performance [Salimans and Kingma, 2016, Huang et al., 2017, Qiao et al., 2020]. Nevertheless, WeightNorm has been one of the most popular normalization methods since its introduction, and as a result current machine learning libraries include built-in implementations of it, e.g., PyTorch 2. .
@@
Parallel to the empirical development of normalization methods, recent years have seen advances in theoretically understanding convergence of gradient descent (GD) and variants for deep learning models [Du et al., 2019, Allen-Zhu et al., 2019, Zou et al., 2020, Nguyen, 2021, Liu et al., 2022, Banerjee et al., 2023]. However, to the best of our knowledge, no prior work has focused on providing formal optimization and generalization guarantees for deep models with WeightNorm. In this paper, our contribution is to study the effect of WeightNorm on both training and generalization from a theoretical perspective. In particular, we focus on smooth networks—deep models with smooth activations—a setting which has become increasingly popular in recent years [Du et al., 2019, Huang and Yau, 2020, Liu et al., 2022, Banerjee et al., 2023].
Training a neural network is a non-convex optimization problem, and establishing optimization guarantees using GD requires analyzing the empirical loss used for training. Particularly, our analysis is based on the second-order Taylor expansion of the empirical loss, for which we require bounds on its Hessian and gradient. We observe that it is possible to bound the empirical loss Hessian by using a bound on the Hessian of the network or predictor. Thus, our first technical contribution (Section 4) is a characterization of the spectral norm of the Hessian of a smooth WeightNorm neural network. Similar to networks without WeightNorm [Liu et al., 2020, Banerjee et al., 2023], our bound decreases with width, but unlike networks without WeightNorm, our bound also decreases with the minimum weight vector norm across all hidden layers. A similar dependence appears when bounding the norm of the empirical loss gradient, which follows from bounding the norm of the predictor gradient (with respect to the network weights). Now, towards establishing the generalization of WeightNorm networks, we obtain a bound on the output of the predictor which does not have any dependence on the minimum weight vector norm. Similarly, the Lipschitz constant of the network does not have such a dependency either. These results are used to subsequently establish a uniform convergence bound based on the Rademacher complexity of WeightNorm networks (Section 6). Remarkably, all of our bounds used for establishing both optimization and generalization guarantees have at most a polynomial dependence on the depth of the network.
Our second technical contribution (Section 5) is to provide optimization guarantees for the training of smooth deep networks with WeightNorm by GD under suitable assumptions. Using our previously derived bounds, we use the restricted strong convexity (RSC) approach recently introduced by Banerjee et al. [2023] to establish sufficient conditions that ensure a decrease on the value of the loss function with respect to its minimum value. In particular, one such condition relies on an inequality (lower bound) on the quantity $\left\lVert\nabla{\cal L}\right\rVert_{2}^{2}/{\cal L}$, where ${\cal L}$ is the empirical loss and $\nabla{\cal L}$ its gradient. This inequality benefits from a larger width as well as a larger minimum weight vector norm–as opposed to another condition for networks without WeightNorm which only benefits from a larger width [Banerjee et al., 2023]. We relate our condition to the so-called restricted Polyak-Lojasiewicz (PL) condition, showing that ours is milder.
@@
Our third technical contribution (Section 6) is to provide generalization guarantees to smooth networks with WeightNorm. We iteratively use WeightNorm properties and our network scaling to obtain a generalization gap which is independent of the network width—our bound is simply ${\cal O}(\sqrt{L}/\sqrt{n})$, where $L$ is the depth and $n$ the sample size. Thus, WeightNorm allows for a natural non-exponential dependence on the depth compared to networks without WeightNorm [Bartlett et al., 2017, Neyshabur et al., 2018, Golowich et al., 2018]. Our generalization guarantee uses a uniform convergence approach based on the Rademacher complexity of WeightNorm networks.
Fourth, we present experiments on the training of WeightNorm networks (Section 7). We present numerical evidence that the convergence rate improves for larger values of (i) the ratio $\left\lVert\nabla{\cal L}\right\rVert_{2}^{2}/{\cal L}$, and (ii) the minimum weight vector norm. According to our theoretical results, larger values of these two quantities strengthen the RSC property of the empirical loss, which then improves the optimization convergence rate.
@@ 2 Related work
WeightNorm was first proposed by Salimans and Kingma [2016], after a year from the introduction of BatchNorm by Ioffe and Szegedy [2015]. LayerNorm was proposed in the same year by Ba et al. [2016], and has also become very popular [Xu et al., 2019]. Extensions and variations of WeightNorm have been introduced since then, for example, centered weight normalization [Huang et al., 2017], orthogonal weight normalization [Huang et al., 2018], and weight standardization [Qiao et al., 2020]. Theoretical analysis of normalization techniques in general has caught an increasing recent interest. We focus mostly on WeightNorm, but mention that BatchNorm has some theoretical work, e.g. [Santurkar et al., 2018, Lian and Liu, 2019]. The work [Arpit et al., 2019] uses a theoretical approach to better come up with initialization schemes tailored for WeightNorm networks with ReLU activations. In contrast, our work focuses in understanding the optimization (training) and generalization of WeightNorm networks with smooth activations. Wu et al. [2020] study the effect of WeightNorm on the convergence of overparameterized least squares problem (which translates to a normalization of the parameters to estimate). We, in contrast, study the non-convex problem of optimizing WeighNorm networks. Dukler et al. [2020] present exponential convergence guarantees on shallow ReLU networks with WeightNorm, whereas we study smooth deep networks and present generalization too. We conclude by remarking that the field of normalization is growing in proposing new normalization schemes and improving their understanding and application—we refer the reader to the recent survey [Huang et al., 2023].
@@
Given the increasingly vast literature, we refer the readers to the surveys [Fan et al., 2021, Bartlett et al., 2021] for an overview of the field of gradient descent training of neural networks. Deep ReLU networks were analyzed by Zou and Gu [2019], Zou et al. [2020], Allen-Zhu et al. [2019], Nguyen [2021], Nguyen et al. [2021], whereas deep networks with smooth activations—our case—were analyzed by Du et al. [2019], Nguyen and Mondelli [2020], Liu et al. [2022]. Regarding the convergence analysis of gradient descent, Du et al. [2019], Allen-Zhu et al. [2019], Zou and Gu [2019], Zou et al. [2020], Liu et al. [2022] used the Neural Tangent Kernel (NTK) approach [Jacot et al., 2018]; whereas Banerjee et al. [2023] used the RSC approach. All training guarantees hold over a neighborhood around the initialization point of the network’s weights.
@@
To the best of our knowledge, the first work in studying the Rademacher complexity for neural networks was [Neyshabur et al., 2015]. Then, the seminal work [Bartlett et al., 2017] provided generalization bounds for neural networks considering Lipschitz activation functions. The bound, obtained using a covering approach, depends exponentially on the depth of the network. The generalization bound was tightened for ReLU activations in [Neyshabur et al., 2018] using a PAC-Bayes approach, but the same exponential dependency persisted. Then, the work [Golowich et al., 2018] avoided some exponential dependence on the depth and established conditions on the norm of the network’s weight matrices in order to obtain generalization bounds independent of depth. Unlike our work, all the cited works do not study WeightNorm networks–-indeed, we show, through a suitable adaptation of such existing results, that WeightNorm allows the generalization bound to depend on $\sqrt{L}$, where $L$ is the depth. Finally, we remark that our generalization bound estimates how well finding a solution to the empirical loss approximates finding a solution to the loss under the real data distribution. A different problem in the literature is inductive bias, i.e., how a specific optimization method, while training the network, may converge to a solution that generalizes well. This was studied for WeightNorm under GD in [Morwani and Ramaswamy, 2022] and under stochastic GD in [Li et al., 2022].
@@ 3 Problem setup: deep learning with WeightNorm
${\left\lVert\bar{W}\right\rVert_{2}}={\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i}\right\rVert_{2}}$. As an example of the notation, let $\{W_{i}^{(l)}\}_{i=1,l=1}^{m,L}$ and $\{{W^{\prime}}_{i}^{(l)}\}_{i=1,l=1}^{m,L}$ be two sets of hidden weights and define $f(W_{i}^{(l)},{W^{\prime}}_{i}^{(l)}):=\alpha W_{i}^{(l)}+\beta{W^{\prime}}_{i}^{(l)}$ for some $\alpha,\beta\in\mathbb{R}$, then $\left\lVert\bar{f}(W,W^{\prime})\right\rVert_{2}=\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\alpha W_{i}^{(l)}+\beta{W^{\prime}}_{i}^{(l)}\right\rVert_{2}$.
@@
Finally, we point out that the original WeightNorm paper by Salimans and Kingma [2016] also has a scalar multiplying the normalized weight vectors $W_{i}^{(l)}/\left\lVert W_{i}^{(l)}\right\rVert_{2}$, $l\in[L]$. Their work considers optimizing over such scalars whereas we consider a fixed scaling for our analysis.
@@ 4 Characterizing WeightNorm bounds on the loss and predictor
The recent works [Liu et al., 2020, Banerjee et al., 2023] analyzed the Hessian spectral norm bound for feedforward networks without WeightNorm. [Liu et al., 2020] presented an exponential dependence on the network’s depth $L$, whereas [Banerjee et al., 2023] avoided exponential dependence upon the choice of the initialization variance for the hidden layers’ weights. We show that WeightNorm eliminates such exponential dependence for at most $L^{5}$ independent from any initialization of the weights, which improves to $L^{3}$ when $\phi(0)=0$. Another important difference is that our bound also depends on the inverse of the minimum weight vector norm—large enough values on the weight vectors decrease the upper bound. Finally, our bound is defined over any value on the weights, and not just over a neighborhood around the initialization point. ∎
@@
Corollary 4.1 shows that the Hessian bound becomes tighter when $\phi(0)=0$ because of an additional $\frac{1}{\sqrt{m}}$ scaling. In contrast, the Hessian bounds by Liu et al. [2020], Banerjee et al. [2023] have the scaling $\frac{1}{\sqrt{m}}$ independent from the value of $\phi(0)$. The reason for this difference is the use of different scale factors: while those two other works introduce an extra scale factor $\frac{1}{\sqrt{m}}$ on the linear output layer, we do not. In our paper, we follow the neural network scaling as done in the seminal work on the Neural Tangent Kernel by Jacot et al. [2018]. ∎
@@
Surprisingly, the Lipschitz constant of the network (7)—unlike networks without WeightNorm—does not explicitly depend on the depth $L$ of the network nor its width $m$. It depends on the radius $\rho_{1}$ which can be set to be a constant ${\cal O}(1)$. This is a direct effect of WeightNorm: when computing $\nabla_{\mathbf{x}}f(\theta;\mathbf{x}_{i})$, we use a chain-rule argument that depends on the product of Jacobians of the output $\alpha^{(l)}$ at layer $l$ with respect to the previous output $\alpha^{(l-1)}$, and we show each Jacobian has at most unit norm due to WeightNorm. Thus, no matter how deep the network is, the final effect this product of Jacobians has on the Lipschitz constant of the network is just a constant one. ∎
@@ 5 Optimization guarantees for WeightNorm
In this section we prove our training guarantees for WeightNorm networks under the square loss.
We first introduce two technical lemmas that will be used for our convergence analysis, defining the restricted strong convexity (RSC) property and a smoothness-like property respectively. These two properties use the bounds derived in Section 4. All proofs are found in Section C of the appendix.
@@
$\begin{split}{\cal L}(\theta^{\prime})&\geq{\cal L}(\theta)+\langle\theta^{\prime}-\theta,\nabla_{\theta}{\cal L}(\theta)\rangle+\frac{\alpha_{\theta,\theta^{\prime}}}{2}\|\theta^{\prime}-\theta\|_{2}^{2}~{},\end{split}$ (12)
@@
$\alpha_{\theta,\theta^{\prime}}=\frac{\kappa^{2}}{2}\frac{\left\lVert\nabla_{\theta}{\cal L}(\theta)\right\rVert_{2}^{2}}{{\cal L}(\theta)}$ (13)
$-{\cal O}\left(\left(1+\frac{\sqrt{L}}{{\left\lVert\bar{W}\right\rVert_{2}}}\right)\frac{(1+\rho_{2})(1+\rho_{1})^{3}L^{3}A(\frac{1}{\sqrt{m}},L^{2},|\phi(0)|)B(1,L,|\phi(0)|\sqrt{m})}{\min\left\{\left\lVert\bar{f}(W^{\prime},W)_{\xi}\right\rVert_{2},\left\lVert\bar{f}(W^{\prime},W)_{\xi}\right\rVert_{2}^{2}\right\}}\right)~{},$
where $f(W^{\prime},W)_{\xi}:=\xi W^{\prime}+(1-\xi)W$ for some $\xi\in[0,1]$, $A(\frac{1}{\sqrt{m}},L^{2},|\phi(0)|):=\frac{1}{\sqrt{m}}+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\}$ and $B(1,L,|\phi(0)|\sqrt{m}):=1+L|\phi(0)|\sqrt{m}$. We say that the empirical loss ${\cal L}$ satisfies the RSC property w.r.t. $(Q^{\theta}_{\kappa}\cap B_{\rho_{2}}^{\operatorname{Euc}}(\theta),\theta)$ whenever $\alpha_{\theta,\theta^{\prime}}>0$.
@@
$\alpha_{\theta,\theta^{\prime}}=\frac{\kappa^{2}}{2}\frac{\left\lVert\nabla_{\theta}{\cal L}(\theta)\right\rVert_{2}^{2}}{{\cal L}(\theta)}$ (14)
$-{\cal O}\left(\left(1+\frac{\sqrt{L}}{\sqrt{m}{\left\lVert\bar{W}\right\rVert_{2}}}\right)\frac{(1+\rho_{2})(1+\rho_{1})^{3}L^{3}}{\sqrt{m}\min\left\{\left\lVert\bar{f}(W^{\prime},W)_{\xi}\right\rVert_{2},\left\lVert\bar{f}(W^{\prime},W)_{\xi}\right\rVert_{2}^{2}\right\}}\right)$
using the same notation as in Lemma 5.1.
@@
Our RSC characterization in the left-hand side of (13) depends on $\frac{\|\nabla_{\theta}\mathcal{L}(\theta)\|_{2}^{2}}{\mathcal{L}(\theta)}$, whereas [Banerjee et al., 2023]—the work that introduced the RSC-based optimization analysis—depends on $\|\frac{1}{n}\sum_{i=1}^{n}\nabla_{\theta}f(\theta,x_{i})\|_{2}^{2}$. Moreover, the minimum weight vector norm appears in the right-hand side since we used the bounds on Section 4. ∎
@@
The RSC condition has the form $\frac{\left\lVert\nabla_{\theta}{\cal L}(\theta)\right\rVert_{2}^{2}}{{\cal L}(\theta)}\geq 2\mu_{\theta,\theta^{\prime}}$ where $\mu_{\theta,\theta^{\prime}}$ depends on the parameters $\theta$ and $\theta^{\prime}$, which are in turn restricted to a specific set. This is a milder condition than the so-called restricted PL condition [Oymak and Soltanolkotabi, 2019, Banerjee et al., 2023], which is satisfied whenever $\mu_{\theta,\theta^{\prime}}$ is a constant independent from both $\theta$ and $\theta^{\prime}$. Our RSC condition becomes even milder when $\phi(0)=0$ (see Corollary 5.1), since now the parameter $\mu_{\theta,\theta^{\prime}}$ has the term $1/\sqrt{m}$ which decreases with the width. ∎
@@
When $\phi(0)=0$, Corollary 5.1 shows that the right-hand side of (13) will have the scaling parameter $\frac{1}{\sqrt{m}}$ due to the Hessian bound (Corollary 4.1). ∎
@@
${\cal L}(\theta^{\prime})\leq{\cal L}(\theta)+\langle\theta^{\prime}-\theta,\nabla_{\theta}{\cal L}(\theta)\rangle+\frac{\beta_{\theta,\theta^{\prime}}}{2}\|\theta^{\prime}-\theta\|_{2}^{2}~{},$ (15)
@@
$\beta_{\theta,\theta^{\prime}}\leq{\cal O}\left(L^{4}A(1,L^{2},|\phi(0)|)(1+\rho_{1})^{2}\left(1+\frac{1}{\min\{\left\lVert\bar{f}(W^{\prime},W)_{\xi}\right\rVert_{2},\left\lVert\bar{f}(W^{\prime},W)_{\xi}\right\rVert_{2}^{2}\}}\right)\right)~{},$ (16)
where $f(W^{\prime},W)_{\xi}:=\xi W^{\prime}+(1-\xi)W$ for some $\xi\in[0,1]$, $A(1,L^{2},|\phi(0)|)=(1+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\}m)^{2}$.
@@
$\beta_{\theta,\theta^{\prime}}\leq{\cal O}\left(1+\frac{L^{3}(1+\rho_{1})^{2}}{\sqrt{m}\min\{\left\lVert\bar{f}(W^{\prime},W)_{\xi}\right\rVert_{2},\left\lVert\bar{f}(W^{\prime},W)_{\xi}\right\rVert_{2}^{2}\}})\right)$ (17)
using the same notation as in Lemma 5.2.
Remark 5.4 (A stronger notion of convexity and smoothness).
If the expressions in equations (12) and (15) hold for any $\theta,\theta^{\prime}\in\mathbb{R}^{p}$ so that $\alpha_{\theta,\theta^{\prime}}\equiv\alpha>0$ and $\beta_{\theta,\theta^{\prime}}\equiv\beta>0$, then we have the definitions of strong convexity and smoothness, respectively—moreover, it immediately follows that $\frac{\alpha}{\beta}<1$. ∎
The following assumptions are useful for establishing our convergence analysis—indeed, as we shortly show, a decrease on the loss value is guaranteed as long as these assumptions are satisfied.
@@
Consider the iterate $\theta_{t}\in\mathbb{R}^{p}$ with $\mathbf{v}_{t}\in B_{\rho_{1}}^{\operatorname{Euc}}(\mathbf{v}_{0})$ and that: (A3.1) gradient descent (GD) update $\theta_{t+1}=\theta_{t}-\eta_{t}\nabla{\cal L}(\theta_{t})$ with learning rate $\eta_{t}>0$ chosen appropriately so that $\mathbf{v}_{t+1}\in B_{\rho_{1}}^{\operatorname{Euc}}(\mathbf{v}_{0})$; (A3.2) $\rho_{2}$ is chosen so that $\theta_{t+1}\in B_{\rho_{2}}^{\operatorname{Euc}}(\theta_{t})$.
The first statement in Assumption 3 ensures that gradient descent has an appropriate learning rate. The second statement defines a ball around the current iterate that should cover the next iterate—in principle, its radius could be defined after defining the learning rate of GD of the first assumption; more details in Remark 5.7.
Finally, we present our main optimization result: a reduction on the empirical loss towards its minimum value (with the last layer in the Euclidean set of radius $\rho_{1}$) using gradient descent. This is proved by using Lemmas 5.1 and 5.2 and an adaptation of [Banerjee et al., 2023, Theorem 5.3]; see Section C.2 of the appendix.